In how many ways can the letters of the word "fish" be arranged if each arrangement is to consist of five letters and repetition is not allowed? if repetition is allowed?
Answer by jim_thompson5910(35256)

You have four letters: f, i, s, h

If you don't allow repeats, then you cannot have a five letter arrangement. You simply do not have enough letters. So there are 0 ways to do this.

If you do allow repeats, then you have 4*4*4*4*4 = 4^5 = 1024 ways to make a five letter "word". One such "word" is fiish. Notice how the 'i' is repeated.


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim